# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ VII Frequency-related applications
The definitions of frequency-related tasks are highly connected, and many algorithms are able to tackle multiple tasks simultaneously.
In the following part, we will first discuss some key issues in handling frequency-related tasks in Section VII-A, and then introduce the existing FA solutions in Section VII-B
VII-A Key issues in frequency-related tasks
@@
Data domain. The data domain (or data universe) is an important setting in frequency-related tasks. The data domain refers to the set of all possible queried data. The easiest setting is the small domain, where the number of possible queried data is small enough so that the algorithm designer can simply enumerate all possible queried data, or design one-hot encoding of them. In the large domain setting, the algorithm designer still has knowledge of all the possible queries, but can no longer apply one-hot encoding or similar techniques, due to the limit of computation and communication resources. In the most difficult infinite domain setting, there are infinite possible queried data. An example of the infinite setting is that the queried data are all possible strings of characters with arbitrary lengths.
@@
VII-B Existing solutions
The majority of FA-based frequency oracle, heavy hitters, and frequent pattern mining solutions adopt LDP for their privacy guarantee, owing to its strong privacy preservation in the untrusted aggregator model. RAPPOR [15] is one of the earliest and most popular frequency oracle solutions with LDP guarantee. It lets the clients encode their local data into bloom filters, so that LDP can be enforced for arbitrary structure of local data. By aggregating the bloom filters, the frequency of the encoded elements can be estimated. [37] (succinct histograms) is another algorithm for frequency oracle and heavy hitter problems. With the help of shared randomness, the succinct histogram algorithm only requests clients to upload one-bit data. LDPMiner [38] is another solution for heavy hitters. Compared to previous solutions, LDPMiner is the first one to resolve the set-value client data, where multiple $q$ can be contained by one $d_{i}$. PSFO [41] proposes its optimization over LDPMiner. While preserving the strengths of LDPMiner, PSFO enables frequent itemset mining tasks with its padding-and-sampling oracle. In [42], two algorithms, TreeHist and Bitstogram, are proposed, to resolve the tasks of frequency oracle, and heavy hitters. In the two algorithms, tailored data structures, shared randomness, and Hadamard transform are carefully designed and utilized. As a result, TreeHist and Bitstrogram manage to optimize the time and space complexity on both the server and client side. In [39], another frequency oracle algorithm based on Hadamard response is proposed that removes the need for shared randomness. In [40], an algorithm is proposed for frequency oracle and heavy hitters. In addition to removing the need for public randomness, [40] only requires one-bit responses. FIML [125] tackles the top-$k$ setting of frequent itemset mining. It first builds a frequency oracle to find out top frequent items, then builds candidate itemsets, and queries their frequencies. OptPrefixTree [107] is a prefix tree-based heavy hitter algorithm. It leverages the prefix tree to handle the infinite data domain.
SFP [82] algorithm is proposed by Apple to tackle the difficult frequent sequence mining problem. It encodes the local client data with count mean sketch so that LDP can be properly enforced. It partially tackles the large domain problem of domain element frequency estimation. It lets clients upload an extra count mean sketch of sequence fragments, where the short fragments can be easily enumerated and frequent short fragments can assemble longer sequences. FedFPM [16] is proposed as a unified FA framework to tackle multiple FPM subproblems, including heavy hitters, frequent itemset mining, and frequent sequence mining. It follows a query-response scheme to estimate the frequency of the candidate patterns and derives Hoeffding’s inequality-based bounds to filter the candidates. FedFPM is able to achieve a better data utility than existing solutions but requires sufficient participating clients.
As LDP schemes request clients to add significant noise to the uploads, researchers propose CDP schemes in FA-based frequency-related tasks. The major advantage of CDP is that it requests clients to add smaller noise on the uploads, or even add no noise by utilizing the sampling mechanism. However, only enforcing CDP sacrifices the formal privacy guarantee, as the FA server is usually considered untrusted. TrieHH [18] is an FA solution to digest string-typed heavy hitters. It iteratively builds up a prefix tree of characters with the interaction between the server and clients. Its major advantage is the ability to satisfy CDP without adding random noise to the outputs so that a better data utility of the mined results can be achieved. TrieHH++ [43] is an expansion of TrieHH. It extends the applications from heavy hitters to quantile estimation and range query. It follows the basic ideas of TrieHH, that realizing CDP via sampling data without adding random noise to the outputs and generating a prefix tree to handle infinite data domain. In addition, it applies Poisson sampling, instead of fixed-size sampling, to hide the sample size.
Some works adopt cryptography tools instead of DP in FA-based frequency-related solutions. Compared to DP, cryptography tools allow deriving accurate results with maximal data utility. However, an adversary still may infer sensitive information from the accurate outputs. STAR [20] is a threshold aggregation (a variation of heavy hitters) solution for web data analysis. Its threshold aggregation enforces $k$-anonymity on the output data, and such functionality is realized by a secret sharing scheme. In [126], a private heavy hitter algorithm is proposed. The solution relies on the incremental distributed point function, a lightweight cryptography tool. It assumes that there exist two servers that collude with neither each other nor any client. In [44], a private heavy hitter solution is proposed that utilizes DDP for its privacy preservation. DDP statistically satisfies CDP, and utilizes cryptography tools to prevent the server from learning any single client upload, so that local privacy can be achieved with CDP-level data utility. In FedWeb [45], another DDP-based solution is proposed. It can handle generic FPM tasks, and proposes tailored design to be applied in Web 3.0 applications.
In [12], an FA-based location heatmap generation solution is proposed, which can be regarded as a frequency oracle solution on geometric items. In [12], DDP, which combines the schemes of cryptography tools and DP, is applied to reduce the noise required for privacy preservation. It iteratively reduces the granularity of queried locations to better utilize the participating clients.
In Table VII, we summarize our surveyed FA-based solutions on frequency-related tasks and list their detailed tasks, privacy model, data domain setting, and extra features.
@@ VII-B Existing solutions
TABLE VII: Summery of FA-based solutions on frequency-related applications.
Reference Task Privacy Data domain Notes
@@ VIII Database operations
As privacy-preserving federated query solutions rely on cryptography tools, especially secure multi-party computation (MPC) protocols, the heavy computation overhead of executing the cryptography tools becomes a significant burden to deploy these federated query solutions into real-world systems. Therefore, an important task of the federated query solutions is to optimize the computation time of executing the federated query and deriving the accurate result. To fulfill that goal, researchers propose various solutions to pursue desirable computation time performance, which utilizes various cryptography tools with different properties and carefully optimizes the computation graph of executing the query to reduce the computation cost.
@@
Federated query solutions usually follow two kinds of architectures. The first architecture is inherited from classical MPC protocols like garbled circuits. In such a fully decentralized architecture, clients directly communicate with each other under the encrypted protocol. After the computation, all the clients can derive the decrypted query results. Another widely applied architecture is the server-client architecture, as is illustrated in Fig. 7. The server is in charge of receiving the SQL statement from the federated query user and composing the computation plan. The clients execute the computation plan to conduct local computation, where client-client communication is usually involved as is required by the encryption tools. The clients upload the encrypted results to the server. Then, the server derives the plaintext SQL result by decrypting the client uploads, and returns the result to the federated query user.
@@
Among the existing federated query solutions, the most ambitious ones are those supporting many types of queries (even the whole SQL standard) within one framework. These solutions are the most powerful FA solutions from the perspective of the number of supported data analytics tasks. SMCQL [46] is a pioneering federated query solution. It utilizes the garbled circuit to encrypt the whole query execution. The utilization of the garbled circuit makes SMCQL have the strong capacity to support the whole SQL standard but introduces significant computation overhead and only supports the computation over two clients. SMCQL also introduces a query planner to optimize the computation graph, so that some computation steps can be performed within one client and do not need SMC execution. Conclave [47] is then proposed to reduce the heavy computation workload of SMCQL. It utilizes the respectively lightweight secret sharing schemes to replace garbled circuit so that the computation is faster; computation among three clients is supported; but some query (e.g., window aggregate) is no longer supported. In addition, Conclave enables clients to optionally annotate some parts (columns) of their data to be non-private. The optional annotation reduces the scale of computation needed for SMC execution and further optimizes the computation time. Senate [50] is another SQL executor over federated data. Compared to SMCQL and Conclave, it changes the semi-honest assumption into a stricter adversarial threat model. It proposes novel SMC decomposition and query planning techniques, letting some computations be executed on a subset of clients. Saqe [49] considers a new direction of optimizing federated query runtime. It removes the restriction that the SQL result must be accurate, and let the federated query derive approximate results. As the execution over a small sampled set of data is enough to derive the query result with a bounded accuracy guarantee, Saqe achieves much better performance by reducing the size of the data to be computed. In addition, the data sampling scheme of Sage realizes an additional DP guarantee of the approximate query result. Orchard [53] is a novel federated query solution that transforms queries into three “zones” that computations on private individual data, aggregated but not noised data, and noised data, respectively. The results of local computation in first “zone” are homomorphically encrypted and sent to the aggregator which performs summation. The aggregated results are decrypted, noised, and announced by a committee later which owns the key. The detailed procedure is the same as its previous work [29] which is introduced in VI-A herein. Orchard can efficiently answer queries at scale as long as there is additive aggregation in them such as k-means, logistic regression, perceptron, and PCA. Arboretum [55] is designed to efficiently answer a wide range of queries in large-scale FA setups with potentially billions of participants. Arboretum’s key strengths include its ability to automatically optimize query plans and distribute computational tasks across participant devices. The system outperforms previous solutions by supporting new types of queries and matches the cost of existing systems that were hand-optimized for specific queries. Arboretum’s approach significantly enhances the scalability and feasibility of executing complex queries in distributed environments while maintaining strong privacy guarantees.
@@
Although the aforementioned general federated query schemes obtain remarkable achievement in computation time reduction and client size scalability, the intrinsic difficulty of the general federated query still burdens its application in many runtime-sensitive and large-scale systems. As an attempt to compromise the types of queries, specific federated queries are proposed. Specific federated query solutions tackle one kind of query so that many tools other than general cryptography solutions can be applied. In other words, researchers only need to provide a cryptography version of a specific algorithm to enable a specific federated query. As a result, the specific federated query has the potential to achieve better performance in computation load and has wide application in scenarios where only a specific kind of query is required. In [48], a federated query solution is proposed to handle the free-connex join-aggregate query (a special kind of query). The authors provide a secure version of the Yannakakis algorithm over two clients, which is much more efficient in runtime compared to the garbled circuit. In [51], the authors consider skyline queries over vertical data federations. Their algorithm design decomposes the execution into local computation and cross-client secure aggregation, where the secure part is reformulated into the private set computation problem, which can be efficiently handled by MPC solutions. Hu-fu [52] tackles SQL queries related to spatial data. The authors notice that various spatial queries, (including kNN, range counting, and range query) can be realized by only three secure operations (summation, comparison, and set union), while other distance-based operations can all be conducted without encryption. Hu-fu then designs effective dedicated algorithms to handle each of the three secure operators and therefore achieves good performance in computation load and gains capacity in supporting up to ten clients. Mycelium [54] is a system that enables processing differentially private queries over large-scale distributed graphs, which are common in scenarios like disease or malware tracking. Mycelium achieves privacy by combining homomorphic encryption , a verifiable secret redistribution scheme (similar to Orchard [53]), and a mix network based on telescoping circuits. It can handle various queries relevant to medical research without compromising individual privacy or learning the graph’s topology.

### v3

@@ VII Frequency-related Applications
In this section, we focus on the second class of data analytics tasks according to our taxonomy: frequency-related applications. Frequency-related applications refer to tasks that involve calculations/estimations on the frequency of particular patterns within the population of clients. After introducing the definitions of frequency-related applications as follows, we summarize some key issues, and then introduce the existing solutions in detail
@@
The definitions of frequency-related tasks are highly connected, and many algorithms can tackle multiple tasks simultaneously.
In the following part, we will first discuss some key issues in handling frequency-related tasks, and then introduce the existing FA solutions.
@@
  • Data domain. The data domain (or data universe) is an important setting in frequency-related tasks. The data domain refers to the set of all possible queried data. The easiest setting is the small domain, where the number of possible queried data is small enough so that the algorithm designer can simply enumerate all possible queried data, or design one-hot encoding of them. In the large domain setting, the algorithm designer still knows all the possible queries, but can no longer apply one-hot encoding or similar techniques, due to the limit of computation and communication resources. In the most difficult infinite domain setting, there are infinite possible queried data. An example of the infinite setting is that the queried data are all possible strings of characters with arbitrary lengths.
@@
We have introduced the key issues in FA applied for frequency-related applications. Then, we introduce the existing FA solutions, sorted by their utilized privatization methodologies.
@@
TABLE VIII: Summery of FA-based solutions on frequency-related applications. (FO: frequency oracle; HH: heavy hitter/frequent item mining; FIsM: frequent itemset mining; FSM: frequent sequence mining)
Reference Task Privacy Data domain Note
@@
Local differential privacy method(s). The majority of FA-based frequency oracle, heavy hitters, and frequent pattern mining solutions adopt LDP for their privacy guarantee, owing to its strong privacy preservation in the untrusted aggregator model. RAPPOR [25] is one of the earliest and most popular frequency oracle solutions with an LDP guarantee. It lets the clients encode their local data into bloom filters so that LDP can be enforced for arbitrary structure of local data. By aggregating the bloom filters, the frequency of the encoded elements can be estimated. [61] (succinct histograms) is another algorithm for frequency oracle and heavy hitter problems. With the help of shared randomness, the succinct histogram algorithm only requests clients to upload one-bit data. LDPMiner [62] is another solution for heavy hitters. Compared to previous solutions, LDPMiner is the first one to resolve the set-value client data, where multiple $q$ can be contained by one $d_{i}$. PSFO [68] proposes its optimization over LDPMiner. While preserving the strengths of LDPMiner, PSFO enables frequent itemset mining tasks with its padding-and-sampling oracle. In [66], two algorithms, TreeHist and Bitstogram, are proposed, to resolve the tasks of frequency oracle, and heavy hitters. In the two algorithms, tailored data structures, shared randomness, and Hadamard transform are carefully designed and utilized. As a result, TreeHist and Bitstrogram manage to optimize the time and space complexity on both the server and client side. In [63], another frequency oracle algorithm based on Hadamard response is proposed that removes the need for shared randomness. In [64], an algorithm is proposed for frequency oracle and heavy hitters. In addition to removing the need for public randomness, [64] only requires one-bit responses. FIML [69] tackles the top-$k$ setting of frequent itemset mining. It first builds a frequency oracle to find out top frequent items, then builds candidate itemsets, and queries their frequencies. OptPrefixTree [67] is a prefix tree-based heavy hitter algorithm. It leverages the prefix tree to handle the infinite data domain.
SFP [65] algorithm is proposed by Apple to tackle the difficult frequent sequence mining problem. It encodes the local client data with the count mean sketch so that LDP can be properly enforced. It partially tackles the large domain problem of domain element frequency estimation. It lets clients upload an extra count mean sketch of sequence fragments, where the short fragments can be easily enumerated and frequent short fragments can assemble longer sequences. FedFPM [26] is proposed as a unified FA framework to tackle multiple FPM subproblems, including heavy hitters, frequent itemset mining, and frequent sequence mining. It follows a query-response scheme to estimate the frequency of the candidate patterns and derives Hoeffding’s inequality-based bounds to filter the candidates. FedFPM can achieve a better data utility than existing solutions but requires sufficient participating clients.
Central differential privacy method(s). As LDP schemes request clients to add significant noise to the uploads, researchers propose CDP schemes in FA-based frequency-related tasks. The major advantage of CDP is that it requests clients to add smaller noise on the uploads, or even add no noise by utilizing the sampling mechanism. However, only enforcing CDP sacrifices the formal privacy guarantee, as the FA server is usually considered untrusted. TrieHH [28] is an FA solution to digest string-typed heavy hitters. It iteratively builds up a prefix tree of characters with the interaction between the server and clients. Its major advantage is the ability to satisfy CDP without adding random noise to the outputs so that a better data utility of the mined results can be achieved. TrieHH++ [70] is an expansion of TrieHH. It extends the applications from heavy hitters to quantile estimation and range query. It follows the basic ideas of TrieHH, that realizing CDP via sampling data without adding random noise to the outputs and generating a prefix tree to handle infinite data domain. In addition, it applies Poisson sampling, instead of fixed-size sampling, to hide the sample size.
Cryptography method(s). Some works adopt cryptography tools instead of DP in FA-based frequency-related solutions. Compared to DP, cryptography tools allow deriving accurate results with maximal data utility. However, an adversary still may infer sensitive information from the accurate outputs. STAR [30] is a threshold aggregation (a variation of heavy hitters) solution for web data analysis. Its threshold aggregation enforces $k$-anonymity on the output data, and such functionality is realized by a secret sharing scheme. In [73], a private heavy hitter algorithm is proposed. The solution relies on the incremental distributed point function, a lightweight cryptography tool. It assumes that there exist two servers that collude with neither each other nor any client.
Distributed differential privacy method(s). In [71], a private heavy hitter solution is proposed that utilizes DDP for its privacy preservation. DDP statistically satisfies CDP, and utilizes cryptography tools to prevent the server from learning any single client upload, so that local privacy can be achieved with CDP-level data utility. In FedWeb [72], another DDP-based solution is proposed. It can handle generic FPM tasks and proposes tailored designs to be applied in Web 3.0 applications. In [17], an FA-based location heatmap generation solution is proposed, which can be regarded as a frequency oracle solution on geometric items. In [17], DDP, which combines the schemes of cryptography tools and DP, is applied to reduce the noise required for privacy preservation. It iteratively reduces the granularity of queried locations to better utilize the participating clients.
In Table VIII, we summarize our surveyed FA-based solutions on frequency-related tasks and list their detailed tasks, privacy model, data domain setting, and extra features.
Lessons learned. FA applied in frequency-related applications focuses on estimating the frequency of any data pattern in the population of clients. Such frequency-related applications can be further classified in many aspects, including dedicated frequency tasks, domain of data, type of data, and type of pattern. These factors determine the difficulty and challenges of the tasks and affect the corresponding FA algorithmic design. Existing solutions utilize LDP, CDP, and cryptography tools for privatization. In addition, as the advanced DP mechanism, DDP is firstly utilized in frequency-related applications and mean computation, because the lightweight encryption tools in DDP are most effective when the insight aggregation phase can be formulated as averaging, where mean computation and frequency-related applications usually follow.
TABLE IX: Summery of federated query solutions.
Reference Query type Privacy Cryptography technique Maximal client
SMCQL [74] General Cryptography Garbled circuit 2
Conclave [75] General Cryptography Secret sharing 3
Senate [78] General Cryptography (adversary model) Garbled circuit 16
SAQE [77] General DDP Secret sharing 2
Orchard [81] General DDP HE $>$1000
Arboretum [83] General DDP HE $>$1000
Secure Yannakakis [76] Free-connex join-aggregate query Cryptography Private set intersection (based on garbled circuit), secret sharing, & oblivious extended permutation 2
Zhang et al. [79] Skyline query Cryptography Private set intersection (based on HE and garbled bloom filter) 10
Hu-fu [80] Spatial query Cryptography Secret sharing 10
Mycelium [82] Graph query DDP HE & onion routing $>$1000
@@ VIII Database Operations
In this section, we focus on the third class of data analytics applications according to our taxonomy: database operations. The FA efforts in this section consider the novel idea of formulating data analytics into SQL queries.
@@
As privacy-preserving federated query solutions rely on cryptography tools, especially MPC protocols, the heavy computation overhead of executing the cryptography tools becomes a significant burden to deploy these federated query solutions into real-world systems. Therefore, an important task of the federated query solutions is to optimize the computation time of executing the federated query and deriving the accurate result. To fulfill that goal, researchers propose various solutions to pursue desirable computation time performance, which utilizes various cryptography tools with different properties and carefully optimizes the computation graph of executing the query to reduce the computation cost.
@@
Federated query solutions usually follow two kinds of architectures. The first architecture is inherited from classical MPC protocols like garbled circuits. In such a fully decentralized architecture, clients directly communicate with each other under the encrypted protocol. After the computation, all the clients can derive the decrypted query results. Another widely applied architecture is the server-client architecture, as is illustrated in Fig. 9. The server is in charge of receiving the SQL statement from the federated query user and composing the computation plan. The clients execute the computation plan to conduct local computation, where client-client communication is usually involved as is required by the encryption tools. The clients upload the encrypted results to the server. Then, the server derives the plaintext SQL result by decrypting the client uploads and returns the result to the federated query user.
Among the existing federated query solutions, the most ambitious ones are those supporting many types of queries (even the whole SQL standard) within one framework. These solutions are the most powerful FA solutions from the perspective of the number of supported data analytics tasks [74, 75, 78, 77, 81, 83]. However, these generate federated query solutions burdens many task-specific applications, and therefore perform bad regarding computation and communication efficiency. Some researchers propose specific federated query solutions as a compromise. Specific federated query solutions tackle one kind of query so that many tools other than general cryptography solutions can be applied [76, 79, 80, 82, 81]. In other words, researchers only need to provide a cryptography version of a specific algorithm to enable a specific federated query. As a result, the specific federated query has the potential to achieve better performance in computation load and has wide application in scenarios where only a specific kind of query is required.
We have concluded the key issues and different types of federated query solutions. Then, we introduce the existing federated query solutions, sorted by the utilized privatization methodologies.
@@
Cryptography method(s). Cryptography tools, like the garbled circuit and secret sharing, are the basis for privacy preservation in federated query solutions. SMCQL [74] is a pioneering federated query solution. It utilizes the garbled circuit to encrypt the whole query execution. The utilization of the garbled circuit makes SMCQL have the strong capacity to support the whole SQL standard but introduces significant computation overhead and only supports the computation over two clients. SMCQL also introduces a query planner to optimize the computation graph, so that some computation steps can be performed within one client and do not need MPC execution. Conclave [75] is then proposed to reduce the heavy computation workload of SMCQL. It utilizes the respectively lightweight secret sharing schemes to replace garbled circuits so that the computation is faster; computation among three clients is supported; but some query (e.g., window aggregate) is no longer supported. In addition, Conclave enables clients to optionally annotate some parts (columns) of their data to be non-private. The optional annotation reduces the scale of computation needed for MPC execution and further optimizes the computation time. Senate [78] is another SQL executor over federated data. Compared to SMCQL and Conclave, it changes the semi-honest assumption into a stricter adversarial threat model. It proposes novel MPC decomposition and query planning techniques, letting some computations be executed on a subset of clients.
@@
Cryptography tool is also a necessity for specific federated query solutions. In [76], a federated query solution is proposed to handle the free-connex join-aggregate query (a special kind of query). The authors provide a secure version of the Yannakakis algorithm over two clients, which is much more efficient in runtime compared to the garbled circuit. In [79], the authors consider skyline queries over vertical data federations. Their algorithm design decomposes the execution into local computation and cross-client secure aggregation, where the secure part is reformulated into the private set computation problem, which can be efficiently handled by MPC solutions. Hu-fu [80] tackles SQL queries related to spatial data. The authors notice that various spatial queries, (including kNN, range counting, and range query) can be realized by only three secure operations (summation, comparison, and set union), while other distance-based operations can all be conducted without encryption. Hu-fu then designs effective dedicated algorithms to handle each of the three secure operators and therefore achieves good performance in computation load and gains capacity in supporting up to ten clients.
Distributed differential privacy method(s). Some researchers design advanced privacy preservation mechanisms for the federated query that combines cryptography tools and CDP, following the idea of DDP, so that an enhanced privacy guarantee can be achieved. SAQE [77] considers a new direction of optimizing federated query runtime. It removes the restriction that the SQL result must be accurate, and let the federated query derive approximate results. As the execution over a small sampled set of data is enough to derive the query result with a bounded accuracy guarantee, SAQE achieves much better performance by reducing the size of the data to be computed. In addition, the data sampling scheme of Sage realizes an additional DP guarantee of the approximate query result. Orchard [81] is a novel federated query solution that transforms queries into three “zones” that computation on private individual data, aggregated but not noised data, and noised data, respectively. The results of the local computation in the first “zone” are homomorphically encrypted and sent to the aggregator which performs summation. The aggregated results are decrypted, noised, and announced by a committee later that owns the key. The detailed procedure is the same as [37]. Orchard can efficiently answer queries at scale as long as there is additive aggregation in them such as k-means, logistic regression, perceptron, and PCA. Arboretum [83] is designed to efficiently answer a wide range of queries in large-scale FA setups with potentially billions of participants. Arboretum’s key strengths include its ability to automatically optimize query plans and distribute computational tasks across participant devices. The system outperforms previous solutions by supporting new types of queries and matching the cost of existing systems that were hand-optimized for specific queries. Arboretum’s approach significantly enhances the scalability and feasibility of executing complex queries in distributed environments while maintaining strong privacy guarantees.
The idea of DDP also applies to specific federated queries. Mycelium [82] is a system that enables processing differentially private queries over large-scale distributed graphs, which are common in scenarios like disease or malware tracking. Mycelium achieves privacy by combining HE, a verifiable secret redistribution scheme (similar to Orchard [81]), and a mixed network based on telescoping circuits. It can handle various queries relevant to medical research without compromising individual privacy or learning the graph’s topology.